Products X, Y, and Z are produced from the same process at a cost of $13,560. Five thousand pounds of raw material yields 3,100 X, 4,100 Y, and 2,600 Z. Selling prices are: X $2 per unit, Y $4 per unit, and Z valueless. The ending inventory of X is 82 units. What is the value of the ending inventory if joint costs are allocated using net realizable value?

Multiple Choice

$89.60.

$98.40.

$100.40.

$80.07.
Answer #1

NRV of X = (3100*2) = 6200

NRV of Y = (4100*4) = 16400

Joint cost is allocated to X on the basis of NRV

Allocation = 13,560 * 6200/(6200+16400)

= 3720

Cost per unit of X = 3720/3100 = 1.20

Codr of ending inventory = 1.20 * 82 units

= 98.40

Option B